please help break this problem down: 6+ sqrt(6)^2-4(1)(10)/2
Answer by rothauserc(4718)

6 +sqrt(6)^2 -4(1)(10)/2 =
6 +sqrt(6)^2 -40/2 =
6 +6 -20 =
-8